# Union of Finitely Generated Congruences on Ground Term Algebra

### v1

@@ 5 Running Examples Continued
$C\langle E,F\rangle=\{\,\{\,\#\,\},\;\{\,f(\#,\#),\;f(\#,g(\#)),\;f(g(\#),\#),\;f(g(\#),g(\#)),\;g(\#),\;g(g(\#))\,\},\;\newline \{\,\$,\;f(\#,\$),\;f(\$,\#),\;f(\$,\$),\;g(\$),\;g(g(\$))\,\}\,\}$.
That is, $C\langle E,F\rangle=\{\,\#/_{\Theta\langle E,F\rangle},\;\$/_{\Theta\langle E,F\rangle},\,f(\#,\#)/_{\Theta\langle E,F\rangle}\,\}$.
$R\langle E,F\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;\$\dot{\rightarrow}\$/_{\Theta\langle E,F\rangle},\;\newline f(\#/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}g(\#)/_{\Theta\langle E,F\rangle},\;f(\#/_{\Theta\langle E,F\rangle},g(\#)/_{\Theta\langle E,F\rangle})\dot{\rightarrow}g(\#)/_{\Theta\langle E,F\rangle},\;\newline f(g(\#)/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}g(\#)/_{\Theta\langle E,F\rangle},\;f(g(\#)/_{\Theta\langle E,F\rangle},g(\#)/_{\Theta\langle E,F\rangle})\dot{\rightarrow}g(\#)/_{\Theta\langle E,F\rangle},\;\newline g(\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}g(\#)/_{\Theta\langle E,F\rangle},\;g(g(\#)/_{\Theta\langle E,F\rangle})\dot{\rightarrow}g(\#)/_{\Theta\langle E,F\rangle},\;\newline f(\#/_{\Theta\langle E,F\rangle},\$/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\$/_{\Theta\langle E,F\rangle},\;f(\$/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\$/_{\Theta\langle E,F\rangle},\;\newline f(\$/_{\Theta\langle E,F\rangle},\$/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\$/_{\Theta\langle E,F\rangle},\;g(\$/_{\Theta\langle E,F\rangle})\dot{\rightarrow}g(\$)/_{\Theta\langle E,F\rangle},\;g(g(\$)/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\$/_{\Theta\langle E,F\rangle}$.
@@
$C\langle F,E\rangle=\{\,\#/_{\Theta\langle F,E\rangle},\;\$/_{\Theta\langle F,E\rangle},\;g(\$)/_{\Theta\langle E,E\rangle},\;f(\#,\$)/_{\Theta\langle E,F\rangle},\;f(\$,\#)/_{\Theta\langle E,F\rangle},\;\newline f(\$,\$)/_{\Theta\langle E,F\rangle},\;g(g(\$))/_{\Theta\langle E,F\rangle}\;\,\}$.
$R\langle F,E\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;\$\dot{\rightarrow}\$/_{\Theta\langle F,E\rangle},\;\newline f(\#/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;f(\#/_{\Theta\langle F,E\rangle},g(\#)/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;\newline f(g(\#)/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;f(g(\#)/_{\Theta\langle F,E\rangle},g(\#)/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;\newline g(\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;g(\$/_{\Theta\langle F,E\rangle})\dot{\rightarrow}g(\$)/_{\Theta\langle F,E\rangle},\;\newline f(\#/_{\Theta\langle F,E\rangle},\$/_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\#,\$)/_{\Theta\langle F,E\rangle},\,f(\$/_{\Theta\langle F,E\rangle},\#_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\$,\#)/_{\Theta\langle F,E\rangle}\newline f(\$/_{\Theta\langle F,E\rangle},\$_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\$,\$)/_{\Theta\langle F,E\rangle},\;g(g(\$)/_{\Theta\langle F,E\rangle})\dot{\rightarrow}g(g(\$))/_{\Theta\langle F,E\rangle}\,\}$.
@@
$C\langle E\cup F,\emptyset\rangle=\ \{\,\#,\;f(\#,\#),\;f(\#,g(\#)),\;f(g(\#),\#),\;f(g(\#),g(\#)),\;g(\#),\;g(g(\#))\,\},\;\newline \{\,\$,\;f(\#,\$),\;f(\$,\#),\;f(\$,\$),\;g(\$),\;g(g(\$))\,\}$.
That is, $C\langle E\cup F,\emptyset\rangle=\{\,\#/_{\Theta\langle E\cup F,\emptyset\rangle},\$/_{\Theta\langle E\cup F,\emptyset\rangle}\,\}$.
$R\langle E\cup F,\emptyset\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;\$\dot{\rightarrow}\$/_{\Theta\langle E\cup F,\emptyset\rangle},\;\newline f(\#/_{\Theta\langle E\cup F,\emptyset\rangle},\#/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;f(\#/_{\Theta\langle E\cup F,\emptyset\rangle},\$/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}\$/_{\Theta\langle E\cup F,\emptyset\rangle},\;\newline f(\$_{\Theta\langle E\cup F,\emptyset\rangle},\#/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}\$/_{\Theta\langle E\cup F,\emptyset\rangle},\;f(\$/_{\Theta\langle E\cup F,\emptyset\rangle},\$/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}\$/_{\Theta\langle E\cup F,\emptyset\rangle},\;\newline g(\$/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}\$/_{\Theta\langle E\cup F,\emptyset\rangle}\,\}$.
@@
$C\langle E,F\rangle=\{\,\{\,\#\,\},\;\{\,\$\,\},\;\{\,f(\#,\#),\;f(\#,\$),\;f(\$,\#),\;f(\$,\$)\,\}\,\}$.
$R\langle E,F\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;\$\dot{\rightarrow}\$/_{\Theta\langle E,F\rangle},\;\newline f(\#/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\#,\#)/_{\Theta\langle E,F\rangle},\;f(\#/_{\Theta\langle E,F\rangle},\$/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\#,\#)/_{\Theta\langle E,F\rangle},\;\newline f(\$/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\#,\#)/_{\Theta\langle E,F\rangle},\;f(\$/_{\Theta\langle E,F\rangle},\$/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\#,\#)/_{\Theta\langle E,F\rangle}\,\},$
$C\langle F,E\rangle=\{\,\{\,\#,\;\$\,\},\;\{\,f(\#,\#),\;f(\#,\$),\;f(\$,\#),\;f(\$,\$)\,\}\,\}$,
$R\langle F,E\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;\$\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;f(\#/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\#,\#)/_{\Theta\langle F,E\rangle}\,\}$,
@@
$C\langle E,F\rangle=\{\,\{\,\#,\;\$\,\},\;\{\,\mathsterling\,\},\;\{\,\flat\,\}\,\}$,
$R\langle E,F\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;\$\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;\mathsterling\dot{\rightarrow}\mathsterling/_{\Theta\langle E,F\rangle},\;\flat\dot{\rightarrow}\flat/_{\Theta\langle E,F\rangle}\,\}$,
$C\langle F,E\rangle=\{\,\{\,\#\,\},\;\{\,\$\,\},\;\{\,\mathsterling,\;\flat\,\}\,\}$,
$R\langle F,E\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;\$\dot{\rightarrow}\$/_{\Theta\langle F,E\rangle},\;\mathsterling\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\flat\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle}\,\}$,
$C\langle E\cup F,\emptyset\rangle=\{\,\{\,\#,\;\$\,\},\;\,\{\,\mathsterling,\;\flat\,\}\,\}$, and
$R\langle E\cup F,\emptyset\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;\$\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;\mathsterling\dot{\rightarrow}\mathsterling/_{\Theta\langle E\cup F,\emptyset\rangle},\;\flat\dot{\rightarrow}\mathsterling/_{\Theta\langle E\cup F,\emptyset\rangle}\,\}$.
@@ 6 The Auxiliary dpwpa for the GTESs $E$ and $F$
From now on throughout the paper, $E$ and $F$ are GTESs over a signature $\Sigma$, $n=size(E)+size(F)$, and $W$ stands for $ST\langle E\cup F\rangle$. We define the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F]\rangle]$ for the GTESs $E$ and $F$. To simplify our notations, each vertex of the dpwpa $AUX[E;F]\rangle$ is a constant $a\in C\langle E\cup F,\emptyset\rangle$, i.e, an equivalence class of $\Theta\langle E\cup F,\emptyset\rangle$. Furthermore, we study $R\langle{E\cup F},\emptyset\rangle$ reachability of a constant $a\in C\langle E\cup F,\emptyset\rangle$ from a proper extension of a constant $b\in C\langle E\cup F,\emptyset\rangle$.
@@
$INC(E\cup F,E)=\{\,(a,b)\mid a\in C\langle E\cup F,\emptyset\rangle,b\in C\langle E,F\rangle,\mbox{ and }b\subseteq a\,\}$,
$NUM(E\cup F,E)=\{\,(a,k)\mid a\in C\langle E\cup F,\emptyset\rangle,k=|\{b\in C\langle E,F\rangle\mid b\subseteq a\,\}|\,\}$,
$INC(E\cup F,F)=\{\,(a,b)\mid a\in C\langle E\cup F,\emptyset\rangle,b\in C\langle F,E\rangle\mbox{ and }b\subseteq a\,\}$, and
$NUM(E\cup F,F)=\{\,(a,k)\mid a\in C\langle E\cup F,\emptyset\rangle,k=|\{b\in C\langle F,E\rangle\mid b\subseteq a\,\}|\,\}$.
Intuitively, $INC(E\cup F,E)$ shows the inclusions between the elements of $C\langle E\cup F,\emptyset\rangle$ and $C\langle E,F\rangle$. $NUM(E\cup F,E)$ consists of all pairs $(a,k)$, where the second component $k$ is the number of all pairs in $INC(E\cup F,E)$, with first component $a$.
Statement 6.59.
@@
By Statement 4.27, $|C\langle E,F\rangle|\leq n$. For each $b\in C\langle E,F\rangle$, there exists at most one $a\in C\langle E\cup F,\emptyset\rangle$ such that $(a,b)\in INC(E\cup F,E)$. Therefore, $|INC(E\cup F,E)|\leq|C\langle E,F\rangle|\leq n$.
@@
Statement 6.60.
1. For every $a\in C\langle E\cup F,\emptyset\rangle$, there exists a unique ordered pair $(a,b)\in INC(E\cup F,E)$ if and only if $a\in C\langle E,F\rangle$. Similarly, for every $a\in C\langle E\cup F,\emptyset\rangle$, there exists a unique ordered pair $(a,b)\in INC(E\cup F,F)$ if and only if $a\in C\langle F,E\rangle$.
2. For every $a\in C\langle E\cup F,\emptyset\rangle$, there exists a unique ordered pair $(a,k)\in NUM(E\cup F,E)$. Here $k=1$ if and only if $a\in C\langle E,F\rangle$. Similarly, for every $a\in C\langle E\cup F,\emptyset\rangle$, there exists a unique ordered pair $(a,\ell)\in NUM(E\cup F,F)$. Here $\ell=1$ if and only if $a\in C\langle F,E\rangle$.
Proof.
By direct inspection of Definition 6.58
@@
Let $E$ and $F$ be GTESs over a signature $\Sigma$, and $a\in C\langle E\cup F,\emptyset\rangle$. We say that GTRS $R\langle E,F\rangle$ keeps up with the GTRS $R\langle E\cup F,\emptyset\rangle$ writing the constant $a\in C\langle E\cup F,\emptyset\rangle$ if
for every rewrite rule of the form $\sigma(a_{1},\ldots,a_{m})\dot{\rightarrow}a\in R\langle{E\cup F},\emptyset\rangle$ with $m\in{\mathbb{N}}$,
[…]
Note that for the constant $b\in C\langle E,F\rangle$ appearing in the definition above, we have $b\subseteq a$. Definition 6.61 also defines the notion that GTRS $R\langle F,E\rangle$ keeps up with the GTRS $R\langle E\cup F,\emptyset\rangle$ writing the constant $a\in C\langle E\cup F,\emptyset\rangle$.
@@
Let $E$ and $F$ be GTESs over a signature $\Sigma$. The auxiliary dpw for the GTESs $E$ and $F$ is $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, where $A[E;F]=BSTEP\langle E\cup F,\emptyset\rangle$. We assign the attributes
@@
For each vertex $a\in$, our decision algorithm computes the values of the above attributes such that the following conditions hold:
@@
    $a.equal_{E}=\mathrm{true}$ if and only if $a\in C\langle E,F\rangle$.
@@
    $a.equal_{F}=\mathrm{true}$ if and only if $a\in C\langle F,E\rangle$.
@@
    $a.keeps_{E}=\mathrm{true}$ if and only if $R\langle E,F\rangle$ keeps up with $R\langle E\cup F,\emptyset\rangle$ writing $a$.
@@
    $a.keeps_{F}=\mathrm{true}$ if and only if $R\langle F,E\rangle$ keeps up with $R\langle E\cup F,\emptyset\rangle$ writing $a$.
[…]
Statement 6.63.
Let $E$ and $F$ be GTESs over a signature $\Sigma$. Then $|A[E;F]|\leq n$.

### v2

@@ 5 Running Examples Continued
[…]
That is, $C\langle E,F\rangle=\{\,\#/_{\Theta\langle E,F\rangle},\;\mathdollar/_{\Theta\langle E,F\rangle},\,f(\#,\#)/_{\Theta\langle E,F\rangle}\,\}$.
$R\langle E,F\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;\mathdollar\dot{\rightarrow}\mathdollar/_{\Theta\langle E,F\rangle},\;\newline f(\#/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}g(\#)/_{\Theta\langle E,F\rangle},\;f(\#/_{\Theta\langle E,F\rangle},g(\#)/_{\Theta\langle E,F\rangle})\dot{\rightarrow}g(\#)/_{\Theta\langle E,F\rangle},\;\newline f(g(\#)/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}g(\#)/_{\Theta\langle E,F\rangle},\;f(g(\#)/_{\Theta\langle E,F\rangle},g(\#)/_{\Theta\langle E,F\rangle})\dot{\rightarrow}g(\#)/_{\Theta\langle E,F\rangle},\;\newline g(\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}g(\#)/_{\Theta\langle E,F\rangle},\;g(g(\#)/_{\Theta\langle E,F\rangle})\dot{\rightarrow}g(\#)/_{\Theta\langle E,F\rangle},\;\newline f(\#/_{\Theta\langle E,F\rangle},\mathdollar/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathdollar/_{\Theta\langle E,F\rangle},\;f(\mathdollar/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathdollar/_{\Theta\langle E,F\rangle},\;\newline f(\mathdollar/_{\Theta\langle E,F\rangle},\mathdollar/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathdollar/_{\Theta\langle E,F\rangle},\;g(\mathdollar/_{\Theta\langle E,F\rangle})\dot{\rightarrow}g(\mathdollar)/_{\Theta\langle E,F\rangle},\;g(g(\mathdollar)/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathdollar/_{\Theta\langle E,F\rangle}$.
@@
$C\langle F,E\rangle=\{\,\{\,\#/_{\Theta\langle F,E\rangle},\;\mathdollar/_{\Theta\langle F,E\rangle},\;g(\mathdollar)/_{\Theta\langle F,E\rangle},\;f(\#,\mathdollar)/_{\Theta\langle F,E\rangle},\;f(\mathdollar,\#)/_{\Theta\langle F,E\rangle},\;\newline f(\mathdollar,\mathdollar)/_{\Theta\langle F,E\rangle},\;g(g(\mathdollar))/_{\Theta\langle F,E\rangle}\,\}\,\}$.
$R\langle F,E\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;\mathdollar\dot{\rightarrow}\mathdollar/_{\Theta\langle F,E\rangle},\;\newline f(\#/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;f(\#/_{\Theta\langle F,E\rangle},g(\#)/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;\newline f(g(\#)/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;f(g(\#)/_{\Theta\langle F,E\rangle},g(\#)/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;\newline g(\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;g(\mathdollar/_{\Theta\langle F,E\rangle})\dot{\rightarrow}g(\mathdollar)/_{\Theta\langle F,E\rangle},\;\newline f(\#/_{\Theta\langle F,E\rangle},\mathdollar/_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\#,\mathdollar)/_{\Theta\langle F,E\rangle},\,f(\mathdollar/_{\Theta\langle F,E\rangle},\#_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\mathdollar,\#)/_{\Theta\langle F,E\rangle}\newline f(\mathdollar/_{\Theta\langle F,E\rangle},\mathdollar_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\mathdollar,\mathdollar)/_{\Theta\langle F,E\rangle},\;g(g(\mathdollar)/_{\Theta\langle F,E\rangle})\dot{\rightarrow}g(g(\mathdollar))/_{\Theta\langle F,E\rangle}\,\}$.
@@
[…]
That is, $C\langle E\cup F,\emptyset\rangle=\{\,\#/_{\Theta\langle E\cup F,\emptyset\rangle},\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}\,\}$.
$R\langle E\cup F,\emptyset\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;\mathdollar\dot{\rightarrow}\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle},\;\newline f(\#/_{\Theta\langle E\cup F,\emptyset\rangle},\#/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;f(\#/_{\Theta\langle E\cup F,\emptyset\rangle},\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle},\;\newline f(\mathdollar_{\Theta\langle E\cup F,\emptyset\rangle},\#/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle},\;f(\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle},\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle},\;\newline g(\#/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;g(\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}\,\}$.
@@
[…]
$R\langle E,F\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;\mathdollar\dot{\rightarrow}\mathdollar/_{\Theta\langle E,F\rangle},\;\newline f(\#/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\#,\#)/_{\Theta\langle E,F\rangle},\;f(\#/_{\Theta\langle E,F\rangle},\mathdollar/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\#,\#)/_{\Theta\langle E,F\rangle},\;\newline f(\mathdollar/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\#,\#)/_{\Theta\langle E,F\rangle},\;f(\mathdollar/_{\Theta\langle E,F\rangle},\mathdollar/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\#,\#)/_{\Theta\langle E,F\rangle}\,\},$
[…]
$R\langle F,E\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;\mathdollar\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;f(\#/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\#,\#)/_{\Theta\langle F,E\rangle}\,\}$,
@@
[…]
$R\langle E,F\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;\mathdollar\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;\mathsterling\dot{\rightarrow}\mathsterling/_{\Theta\langle E,F\rangle},\;\flat\dot{\rightarrow}\flat/_{\Theta\langle E,F\rangle}\,\}$,
[…]
$R\langle F,E\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;\mathdollar\dot{\rightarrow}\mathdollar/_{\Theta\langle F,E\rangle},\;\mathsterling\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\flat\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle}\,\}$,
[…]
$R\langle E\cup F,\emptyset\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;\mathdollar\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;\mathsterling\dot{\rightarrow}\mathsterling/_{\Theta\langle E\cup F,\emptyset\rangle},\;\flat\dot{\rightarrow}\mathsterling/_{\Theta\langle E\cup F,\emptyset\rangle}\,\}$.
@@ 6 The Auxiliary dpwpa for the GTESs $E$ and $F$
Recall that $E$ and $F$ are GTESs over a signature $\Sigma$, $n=\mathrm{size}(E)+\mathrm{size}(F)$, and $W$ stands for $ST\langle E\cup F\rangle$. We define the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ for the GTESs $E$ and $F$. To simplify our notations, each vertex of the dpwpa $AUX[E;F]$ is a constant ${b}\in C\langle E\cup F,\emptyset\rangle$, i.e, an equivalence class of $\Theta\langle E\cup F,\emptyset\rangle$. Furthermore, we study the $R\langle{E\cup F},\emptyset\rangle$ reachability of a constant ${b}\in C\langle E\cup F,\emptyset\rangle$ from a proper extension of a constant ${c}\in C\langle E\cup F,\emptyset\rangle$.
@@
$INC(E\cup F,E)=\{\,({b},{c})\mid{b}\in C\langle E\cup F,\emptyset\rangle,{c}\in C\langle E,F\rangle,\mbox{ and }{c}\subseteq{b}\,\}$,
$NUM(E\cup F,E)=\{\,({b},k)\mid{b}\in C\langle E\cup F,\emptyset\rangle,k=|\{{c}\in C\langle E,F\rangle\mid{c}\subseteq{b}\,\}|\,\}$,
$INC(E\cup F,F)=\{\,({b},{c})\mid{b}\in C\langle E\cup F,\emptyset\rangle,{c}\in C\langle F,E\rangle\mbox{ and }{c}\subseteq{b}\,\}$, and
$NUM(E\cup F,F)=\{\,({b},k)\mid{b}\in C\langle E\cup F,\emptyset\rangle,k=|\{{c}\in C\langle F,E\rangle\mid{c}\subseteq{b}\,\}|\,\}$.
Intuitively, $INC(E\cup F,E)$ shows the inclusions between the elements of $C\langle E\cup F,\emptyset\rangle$ and $C\langle E,F\rangle$. $NUM(E\cup F,E)$ consists of all pairs (b,k)({b},k), where the second component $k$ is the number of all pairs in $INC(E\cup F,E)$, with first component b{b}. Similar statements hold true for $INC(E\cup F,F)$ and $NUM(E\cup F,F)$ as well.
Statement 6.60.
@@
By Statement 4.28, $|C\langle E,F\rangle|\leq n$. For each ${c}\in C\langle E,F\rangle$, there exists at most one ${b}\in C\langle E\cup F,\emptyset\rangle$ such that $({b},{c})\in INC(E\cup F,E)$. Therefore, $|INC(E\cup F,E)|\leq|C\langle E,F\rangle|\leq n$.
@@
We get the following statement by direct inspection of Definition 6.59.
Statement 6.61.
1. For every ${b}\in C\langle E\cup F,\emptyset\rangle$, there exists a unique ordered pair $({b},{c})\in INC(E\cup F,E)$ if and only if ${b}\in C\langle E,F\rangle$. Similarly, for every ${b}\in C\langle E\cup F,\emptyset\rangle$, there exists a unique ordered pair $({b},{c})\in INC(E\cup F,F)$ if and only if ${b}\in C\langle F,E\rangle$.
2. For every ${b}\in C\langle E\cup F,\emptyset\rangle$, there exists a unique ordered pair $({b},k)\in NUM(E\cup F,E)$. Here $k=1$ if and only if ${b}\in C\langle E,F\rangle$. Similarly, for every ${b}\in C\langle E\cup F,\emptyset\rangle$, there exists a unique ordered pair $({b},\ell)\in NUM(E\cup F,F)$. Here $\ell=1$ if and only if ${b}\in C\langle F,E\rangle$.
@@
Let $E$ and $F$ be GTESs over a signature $\Sigma$, and ${b}\in C\langle E\cup F,\emptyset\rangle$. We say that GTRS $R\langle E,F\rangle$ keeps up with the GTRS $R\langle E\cup F,\emptyset\rangle$ writing the constant ${b}\in C\langle E\cup F,\emptyset\rangle$ if
for every rewrite rule of the form $\sigma({b}_{1},\ldots,{b}_{m})\dot{\rightarrow}{b}\in R\langle{E\cup F},\emptyset\rangle$ with
$m\in{\mathbb{N}}$, $\sigma\in\Sigma^{(m)}$, and ${b}_{1},\ldots,{b}_{m}\in C\langle E\cup F,\emptyset\rangle$,
for every ${c}_{1},\ldots,{c}_{m}\in C\langle E,F\rangle$ with ${c}_{i}\subseteq{b}_{i}$ for every $i\in\{\,1,\ldots,m\,\}$,
there exists a rewrite rule in $R\langle E,F\rangle$ of the form $\sigma({c}_{1},\ldots,{c}_{m})\dot{\rightarrow}{c}$ for some ${c}\in C\langle E,F\rangle$.
Note that for the constant ${c}\in C\langle E,F\rangle$ appearing in the definition above, we have ${c}\subseteq{b}$. Definition 6.62 also defines the notion that GTRS $R\langle F,E\rangle$ keeps up with the GTRS $R\langle E\cup F,\emptyset\rangle$ writing the constant ${b}\in C\langle E\cup F,\emptyset\rangle$.
@@
Let $E$ and $F$ be GTESs over a signature $\Sigma$. The auxiliary dpwpa for the GTESs $E$ and $F$ is $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, where $A[E;F]=BSTEP\langle E\cup F,\emptyset\rangle$. We assign the attributes
@@
For each vertex ${b}\in C\langle E\cup F,\emptyset\rangle$, our decision algorithm computes the values of the above attributes such that the following conditions hold:
@@
    ${b}.equal_{E}=\mathrm{true}$ if and only if ${b}\in C\langle E,F\rangle$.
@@
    ${b}.equal_{F}=\mathrm{true}$ if and only if ${b}\in C\langle F,E\rangle$.
@@
    ${b}.keeps_{E}=\mathrm{true}$ if and only if $R\langle E,F\rangle$ keeps up with $R\langle E\cup F,\emptyset\rangle$ writing b{b}.
@@
    ${b}.keeps_{F}=\mathrm{true}$ if and only if $R\langle F,E\rangle$ keeps up with $R\langle E\cup F,\emptyset\rangle$ writing b{b}.
Recall that for each vertex ${b}\in C\langle E\cup F,\emptyset\rangle$, list Adj(b)Adj({b}) consists of all vertices ${c}\in C\langle E\cup F,\emptyset\rangle$ such that there exists a rewrite rule $\sigma({b}_{1},\ldots,{b}_{m})\dot{\rightarrow}{b}$ in $R\langle E\cup F,\emptyset\rangle$ with $\sigma\in\Sigma_{m}$, $m\geq 1$, and ${c}={b}_{i}$ for some $i\in\{\,1,\ldots,m\,\}$.
Statement 6.64.
Let $E$ and $F$ be GTESs over a signature $\Sigma$. Then $|A[E;F]|\leq 2n$.